In a lottery game the six winning numbers are chosen at random from
the set f {1,2,3 ..... 36}. Your entry consists of a similar choice. Find the
probability of:
(a) a win (i.e. your entry matches the winning numbers)
(b) a match four (i.e. exactly four of your entries have been chosen as
winning numbers)

Answer by stanbon(75887)

In a lottery game the six winning numbers are chosen at random from
the set f {1,2,3 ..... 36}. Your entry consists of a similar choice. Find the
probability of:
(a) a win (i.e. your entry matches the winning numbers)
# of sets of 6 numbers = 36C6 = 1947792
# of winniing sets of 6 numbers = 1
P(winning) = 1/1947792
Suggestion: Save your money
------------------------------------------
(b) a match four (i.e. exactly four of your entries have been chosen as
winning numbers)
# of sets of 4 of the winning numbers = 6C4 = 6C2 = 6*5/1*2 = 15
P(match of four) = 15/1947792
------------------------------------------
Cheers,
Stan H.